# Revisiting Stochastic Approximation and Stochastic Gradient Descent

### v2

@@ 3 Proofs
Of Theorem 8 We start by noting that under conditions (H1) as well as the case $\alpha=2$ under (H5), the conditions of Theorem 6 are satisfied with $A_{n,i}^{c}=\emptyset$ (the empty set). Condition (20) holds trivially and (24) holds because of the condition on $\{W_{n}\}$ that $E(W_{n})=0$ under (H1) and $\{W_{n}\}$ being a martingale difference sequence under $(H5)$. Similarly, (22) holds in view of (25) under (H1) and (30) under (H5), in the case $\alpha=2$. Gencs GSLLN holds.
@@
in view of the assumption (30). Thus (21) holds. Coming to the third condition, using $a^{\alpha}\geq a^{2}$ when $0<a\leq 1$ and $21<\alpha<$ 2, we have
@@
Coming to the (22), we have (recall that $21\leq\alpha<$ 2)
@@
it follows that $\sum_{i}{\mathcal{P}}(A_{n+1,i})<\infty$ and hence (20) holds. Since $\{W_{n}:n\geq 1\}$ are i.i.d. with symmetric distribution (that is, $W_{1}$ and $-W_{1}$ have the same distribtuion), it follows that $W_{n,i}\mbox{{ \sf 1}}_{A_{n+1,i}}$ has symmetric distribution for each $i$; hence (21) holds trivially. Coming to (22), it remains to show that

### v3

@@ 3 Proofs
Of Theorem 8. We start by noting that under conditions (H1) as well as the case $\alpha=2$ under (H5), the conditions of Theorem 6 are satisfied with $A_{n,i}^{c}=\emptyset$ (the empty set). Condition (20) holds trivially and (24) holds because of the condition on $\{W_{n}\}$ that $E(W_{n})=0$ under (H1) and $\{W_{n}\}$ being a martingale difference sequence under $(H5)$. Similarly, (22) holds in view of (25) under (H1) and (30) under (H5), in the case $\alpha=2$. Hence GSLLN holds.
@@
in view of the assumption (30). Thus (21) holds. Coming to the third condition, using $a^{\alpha}\geq a^{2}$ when $0<a\leq 1$ and $1<\alpha<2$, we have
@@
Coming to the (22), we have (recall that $1\leq\alpha<2$)
@@
it follows that $\sum_{i}{\mathcal{P}}(A_{n+1,i})<\infty$ and hence (20) holds. Since $\{W_{n}:n\geq 1\}$ are i.i.d. with symmetric distribution (that is, $W_{1}$ and $W_{1}$ have the same distribtuion), it follows that $W_{n,i}\mbox{{ \sf 1}}_{A_{n+1,i}}$ has symmetric distribution for each $i$; hence (21) holds trivially. Coming to (22), it remains to show that